OpenStudy (anonymous):

K is the midpoint of AB if AK =3x-4, and KB= 1/2x+6 find the length of AB

OpenStudy (goformit100):

Is there any figure or diagram for this question ?

OpenStudy (mathstudent55):

 Set the length of AK equal to the length of KB and solve for x. Then substitute x into 3x - 4 to find the length of AK. Then multiply it by 2 to find the length of AB.

OpenStudy (anonymous):

can you help me do that, ive been out of school for a while im not quite sure how to do any of this

OpenStudy (anonymous):

Well if you know that k is the midpoint then you know that AK is equal to KB. Why because taht is teh defintion of midpoint; A midpoint divides something in equal pieces. So set AK=KB and solve for x So 3x-4=1/2x+6... Once you solve for x. Plug in x for AK+KB cause AK+KB=AB and that is what you want. Does that make sense?

OpenStudy (anonymous):

yes i just dont know how to solve for x

OpenStudy (anonymous):

Ok wellc solving for x is basic algebra. Your goal is to get the x by itself on one side and the numbers on the other side.. So what we do is 3x-4=1/2x+6 (I want to get x's on left side and numbers on left side) So add 4 to both sides 3x=1/2x+10 Then subtract 1/2x to both sides 1.5x=10 N0w to get rid of the 1.5 Divide by 1.5 to both sides... x=6 2/3

OpenStudy (anonymous):

so the equation would be 3(6 2/3)-4=1/2(6 2/3)+6?

OpenStudy (anonymous):

Plug in x for AK+KB cause AK+KB=AB and that is what you want. Not AK=KB we want to add them not set them equal to each other

OpenStudy (anonymous):

ok so them added together is 25.1 is that my answer?

OpenStudy (mathstudent55):

@Gerardo_cast23 Ok wellc solving for x is basic algebra. Your goal is to get the x by itself on one side and the numbers on the other side.. So what we do is 3x-4=1/2x+6 (I want to get x's on left side and numbers on left side) So add 4 to both sides 3x=1/2x+10 Then subtract 1/2x to both sides 1.5x=10 N0w to get rid of the 1.5 Divide by 1.5 to both sides... ^^^^^^^^^^^^^^^^^^^^^^^ Careful with your arithmetic: 3x - (1/2)x = 3x - 0.5x = 2.5x, not 1.5 x. x = 4, not x = 6 2/3 x=6 2/3

OpenStudy (anonymous):

Oh oops sorry>.<

OpenStudy (mathstudent55):

3x - 4 = (1/2)x + 6 Replace 1/2 with 0.5 3x - 4 = 0.5x + 6 Subtract 0.5x from both sides: 2.5x - 4 = 6 Add 4 to both sides: 2.5x = 10 Divide both sides by 2.5 x = 4

OpenStudy (mathstudent55):

Now that we know x = 4, we can find AK: AK = 3x - 4 AK = 3(4) - 4 AK = 12 - 4 AK = 8 Since AB = 2(AK) AB = 2(8) AB = 16

OpenStudy (anonymous):

so my answer is 16?

OpenStudy (mathstudent55):

Correct.

OpenStudy (anonymous):

Thank you for checking/finding my mistake

OpenStudy (anonymous):

ok thank you guys!

OpenStudy (mathstudent55):

You're welcome. Good explanation, Gerardo. Anyone can make a silly mistake. I've made plenty of them. :-)